What two numbers is the sum of 20 and when the second number is subtracted by the first number, the difference is 14?
Answer by rm29924(97)

let x = the first number
let y = the second number
the sum of the numbers is 20
so x+y=20
when the second number is subtracted by the first number the difference is 14
so x-y=14
solve for x in the second equation by adding y to both sides
x=14+y
now plug that value for x into the first equation
14+y+y=20
14+2y=20
subtract 14 from both sides
2y=6
divide 2 from both sides to get the variable y by itself
y=3
x=14+y so x=14+3=17
answer: x=17 and y=3